A study done by researchers at a university concluded that 30​% of all student athletes in this country have been subjected to some form of hazing. The study is based on responses from 1,800 athletes. What are the margin of error and​ 95% confidence interval for the​ study?
Answer #1

Given that, proportion ( p ) = 0.30

sample size ( n ) = 1800

A 95% confidence level has significance level (alpha) = 0.05 and critical value is, af2-1.96

We want to find, the margin of error ( E ),

z,/2 * V/p*(1-P) = 196 * V/0.30 * (1 00.30) = 0.021170 E = Za/2 * 1800

Therefore, the margin of error is 0.0212
Answer #2
A research calculated the percentages of students who are female at a sample of colleges and universities. The unit of analysis in this study was
source: إحصاء
answered by: Aaaa